# Unsupervised Representation Learning for 3D Mesh Parameterization with Semantic and Visibility Objectives

### v2

@@ 3.4 Visibility-Aware UV Parameterization
To address this, we introduce a visibility-aware objective that biases seam placement toward surface areas with low viewer exposure. As a proxy for visibility, we leverage ambient occlusion (AO) Zhukov et al. (1998), a standard geometric measure of how much of the local hemisphere around a surface point is blocked by nearby geometry Landis (2002); Christensen (2002); Bavoil & Sainz (2008); McGuire (2010); Laine & Karras (2010); Méndez-Feliu & Sbert (2009). Our training procedure consists of four steps: (i) compute per-vertex AO values on the input mesh, (ii) apply the base UV-parameterization backbone (Sec. 3.2) to obtain candidate UV islands, (iii) extract UV boundary points corresponding to cutting seams, and (iv) compute and minimize the AO-weighted average over the associated 3D vertices. This objective guides the backbone to produce UV islands whose seams coincide with regions of low exposure, thereby relocating cuts to less visible areas of the 3D surface and reducing perceptible seam artifacts after texturing and rendering. We now detail each step in turn.
@@
By this convention, $AO(p)=1$ indicates a fully exposed point (no occlusion) and $AO(p)=0$ indicates a fully occluded point. In practice AO is estimated by Monte Carlo hemisphere sampling Veach (1998) or by high-quality offline ray sampling Pharr et al. (2023) for evaluation. In our implementation, we compute ambient occlusion using the libigl geometry processing library Jacobson et al. (2018), leveraging its mesh-based AO utilities for efficient per-vertex AO estimation.
Learning Mesh Parameterization. Once per-vertex AO values are computed, we generate the UV parameterization using the base backbone described in Sec. 3.2 and Appendix. A.2 to obtain a potential UV parameterization candidate.
@@
Method Visibility $\downarrow$ Conformality $\uparrow$ Equiareality $\uparrow$ Time (s) $\downarrow$ OptCuts Li et al. (2018) 0.7855 0.9341 0.8934 240 FlexPara (Single-Chart) Zhao et al. (2025) 0.8604 0.9097 0.6759 2 Ours - Visibility-Aware Param 0.6065 0.9175 0.6093 2 Ours - Semantic+Visibility-Aware Param 0.6534 0.9153 0.6369 15
Visibility-Aware Seam Loss. After identifying soft seam memberships $s_{i}\in[0,1]$ for each vertex $v_{i}$, we encourage seams to lie in less visible regions of the surface. As a visibility proxy, we use per-vertex ambient occlusion values $AO_{i}\in[0,1]$, computed in the first step, where 0 indicates fully occluded and 11 indicates fully visible. The loss is defined as the weighted average of occlusion values, using soft seam memberships as weights:
@@
Figure 4: Qualitative results for visibility-aware seam placement and UV parameterization on three representative meshes. For each mesh, the top row shows per-vertex ambient occlusion (yellow = exposed, purple = occluded). Beneath are the visualizations of cutting seams (red) from our method, FlexPara Zhao et al. (2025), and OptCuts Li et al. (2018) (top to bottom). Our method places a larger fraction of seam geometry in less-exposed regions, reducing the likelihood of visible seam artifacts under typical viewpoints.
@@ 4 Experiments and Results
We evaluate our unsupervised semantic- and visibility-aware UV parameterization pipelines on a diverse collection of meshes and compare against several baselines. We provide both qualitative and quantitative comparisons to: FlexPara Zhao et al. (2025) (recent state-of-the-art global parameterization network), Autodesk Maya Autodesk, Inc. (2025) (industry-grade commercial software), Blender Blender Foundation (2025) (popular open-source modeling tool), and xatlas jpcy (2025) (widely-used C++ library for generating unique texture coordinates suitable for lightmap baking and texture painting).
Figure 5: Checkerboard texturing comparison using UV parameterizations produced by our visibility-aware method, FlexPara Zhao et al. (2025), and OptCuts Li et al. (2018). Each row shows rendered views of different meshes textured with a checkerboard and a magnified inset of a visually important region near seams (red circles). Because our method steers seams toward occluded regions, the checkerboard pattern appears substantially more continuous from typical camera viewpoints. By contrast, baselines exhibit visible seam artifacts in the zoomed-in insets.
Figs. 1,  3,  4, and  5 show representative qualitative results, and Tables. 1 and  3 presents quantitative results of our semantic-aware and visibility-aware pipelines against baselines (additional results are provided in Appendix A.3). In the following, we detail each experiment and its corresponding results.
@@ 4.1 Qualitative Evaluation
Comparative Evaluation of Semantic-Aware UV Parameterization. To evaluate the performance of the proposed semantic-aware UV parameterization pipeline, we compare our method against FlexPara Zhao et al. (2025), Autodesk Maya’s Automatic UV generator Autodesk, Inc. (2025), Blender’s SmartUV Blender Foundation (2025), and xatlas jpcy (2025).Fig. 3 shows a representative comparison on a single Bunny mesh; additional results on more diverse and complex meshes are provided in Figs. 12 and 13 in Appendix A.3. In Fig. 3 and Fig. 12, each row corresponds to one method. Semantic parts are color-coded, and columns show the rendered 3D mesh from multiple camera viewpoints. The rightmost column depicts the unwrapped UV coordinates, where UV islands are color-coded consistently with their corresponding semantic 3D parts in the rendered views. As shown, our method successfully identifies semantic parts of the input mesh, parameterizes them, and maps them into UV space with low distortion. This leads to cleaner, more interpretable charts that directly benefit downstream tasks such as texture painting and transfer. By contrast, the baseline methods fail to consistently capture and parameterize semantic parts of the 3D mesh.
Comparative Evaluation of Visibility-Aware UV Parameterization. To evaluate the performance of the proposed visibility-aware UV parameterization pipeline, we compare our method against FlexPara Zhao et al. (2025) and OptCuts Li et al. (2018). Figs. 4 and  9 present representative comparisons on different meshes. Additional examples appear in Figs. 5,  10, and 11 in Appendix A.3. In Fig. 4 the top row visualizes per-vertex ambient occlusion (AO): lighter tones (e.g., yellow) indicate high exposure (more visible) while darker tones (e.g., purple) indicate low exposure (more occluded). A desirable parameterization therefore places cutting seams on darker, less-visible regions to reduce perceptible seam artifacts after texturing and rendering. The rows beneath show UV parameterizations produced by our method, FlexPara, and OptCuts (in that order). Qualitatively, our method allocates a larger fraction of seam geometry to low-exposure regions compared to the baselines. Table 1 (and Fig. 8 in Appendix A.3) quantitatively supports this observation: the report the mean AO values of seam vertices for each method, where lower scores indicate seams located in less-visible (more occluded) areas. As shown, our method consistently attains lower mean seam-AO, confirming that the visibility-aware objective successfully biases seam placement toward less-exposed surface regions. Also, Fig. 5 illustrates checkerboard texturing on different mesh objects using parameterizations from our method and from baselines. As shown, our method produces noticeably more seamless texture patterns when rendered from visible viewpoints, whereas FlexPara and OptCuts exhibit prominent cutting seams (highlighted in the zoomed-in insets of Fig. 5).
@@ 4.2 Quantitative Evaluation
User Study. To evaluate semantic- and visibility-awareness of the proposed method, we conducted a user-preference study with 115 participants. We distributed the questionnaire to two groups: experts and general participants. Of the 115 participants, 70 are general participants (including graduate students with computer science and engineering backgrounds) and 45 are experts: 31 software engineers, 3 project managers, 2 product owners, 5 UV/layout artists, and 4 modeling artists, all working in the Media and Entertainment industry for film and games.
@@
Table 2: To evaluate semantic- and visibility-awareness of the proposed method, we conducted a user‐study with 45 expert participants performing 11 comparisons between textured 3D shapes and UV parameterizations produced by our method and baselines. We report the percentage of participant preferences for each method. Our proposed method is strongly preferred by the expert users over the baselines.
@@
Method Expert User Preference Percentage Visibility-Awareness Evaluation OptCuts Li et al. (2018) 1.48 % FlexPara (single-Chart) Zhao et al. (2025) 4.81 % Ours - Visibility-Aware Param 93.70 % Semantic-Awareness Evaluation xatlas jpcy (2025) 2.66 % Blender Blender Foundation (2025) 5.33 % Autodesk Maya Autodesk, Inc. (2025) 8.44 % FlexPara (multi-Chart) Zhao et al. (2025) 3.33 % Ours - Semantic-Aware Param 80.22 %
@@
Each participant completed 11 comparisons between textured 3D shapes and UV parameterizations produced by our method, FlexPara Zhao et al. (2025), OptCuts Li et al. (2018), Autodesk Maya Autodesk, Inc. (2025), Blender Blender Foundation (2025), and xatlas jpcy (2025). For each comparison, participants were rated each result according to three visual criteria: (i) texture-pattern smoothness/continuity, (ii) semantic alignment (how well colors correspond to meaningful parts), and (iii) seam visibility (how well seams are placed in occluded/less-obvious regions). The semantic-aware and visibility-aware evaluations comprised 5 and 6 questions, respectively, for a total of 11 comparisons per participant. Table 2 reports the percentage of expert participant preferences for each method. As the table shows, expert users, particularly UV, layout, and modeling artists, strongly prefer our proposed method over the baselines. Examples of questions in our user‐study questionnaire are shown in Fig. 16. Also, Table 10 in Appendix A.3 reports the percentage of general participant preferences for each method. As shown in Table 10, general users also strongly prefer our proposed method over the baselines.
@@
Semantic- and Visibility-Awareness Evaluation. We also quantitatively evaluate both the semantic-aware and visibility-aware components of our pipeline on a collection of representative meshes. For the semantic-aware pipeline we report (i) Hamming distance (lower is better), the fraction of vertices whose labels disagree after optimal label matching to the reference (we use SAMesh Tang et al. (2024) outputs as a ground-truth); and (ii) Rand Index (RI) (higher is better), a pairwise agreement measure that counts how often vertex pairs are grouped or separated identically between the two labelings. For the visibility-aware pipeline we report Mean seam AO (lower is better in our AO convention, where 0 = occluded and 1 = exposed), the average ambient-occlusion value over seam vertices, which indicates how well seams are placed in occluded regions. Tables 1 and  3 summarize these metrics and compare our method against the baselines described in Sec. 4. Further discussion of the results in these two tables are provided in Appendix A.3.
@@
Geometry-Preservation Evaluation. In addition to task-specific metrics mentioned above, in Tables 1 and  3, we evaluate geometric quality for all methods using conformality (angle-preservation) and equiareality (area-preservation) (higher is better) to ensure the parameterizations remain faithful enough to the original surface. The proposed semantic-aware and visibility-aware UV parameterization frameworks produce an increase in standard distortion metrics including angle-distortion and area-distortion. This is a trade-off which is expected when optimizing multiple objectives during the learning process. Specifically, our optimization introduces explicit semantic-alignment and visibility-awareness terms that encourage semantically consistent charts and reduce conspicuous seams. However, existing geometry-preserving methods (e.g., Autodesk Maya Autodesk, Inc. (2025), Blender Blender Foundation (2025), xatlas jpcy (2025), and FlexPara Zhao et al. (2025)) focus only on traditional distortion objectives and therefore score better on those numbers, but they do not provide the semantic- and visibility-related guarantees produced by our method.
@@
Method Semantic Awareness Conformality $\uparrow$ Equiareality $\uparrow$ Inference Time (sec) $\downarrow$ Hamming Distance $\downarrow$ Rand Index $\uparrow$ xatlas jpcy (2025) 0.8896 0.7023 0.9792 0.9341 12 Blender Blender Foundation (2025) 0.8634 0.7173 0.9289 0.9023 << 1 Maya Autodesk, Inc. (2025) 0.8615 0.7125 0.9272 0.8746 << 1 FlexPara (Multi-Chart) Zhao et al. (2025) 0.5980 0.6902 0.9592 0.7606 2 Ours - Semantic-Aware Param 0.3188 0.8151 0.9123 0.6707 15 Ours - Semantic+Visibility-Aware Param 0.3212 0.8087 0.9153 0.6369 15
@@
Ablation Studies. To understand which design choices drive our results, we perform a set of controlled ablations and summarize the key findings below (additional qualitative examples are in Fig. 14 and Fig. 15 in Appendix A.3). (i) Partitioning strategy (ShDF vs. SAMesh). We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively both partitioners produce part decompositions that align with geometric structure across simple and complex models, resulting compact, low-distortion per-part UV charts (see the appendix Fig. 14 and Fig. 15).
Figure 6: Ablation: Quantitative distribution of seam vertices ambient occlusion for meshes parameterized by using our proposed method, FlexPara Zhao et al. (2025), and OptCuts Li et al. (2018).
Quantitatively (Table 6 in Appendix. A.3), using ShDF in our pipeline yields UV parameterizations that are more conformal (better angle preservation) and more equiareal (better area preservation) than those produced when using SAMesh. (ii) Visibility loss (with vs. without). We ablate the proposed visibility-aware objective (Sec. 3.4) by training identical models with and without $\mathcal{L}_{\text{vis}}$. The visibility-aware objective consistently lowers the mean AO of seam vertices (Fig 6), i.e., it moves seams to regions of lower exposure (lower AO values). Qualitatively, models trained with the AO loss produce seams that are less salient in rendered images from canonical viewpoints (see Fig. 5). Importantly, as presented in Table. 1, adding the visibility term does not noticeably degrade geometric distortion metrics, conformality (angle-preservation) and equiareality area-preservation). The ablations show that: (i) the ShDF partitioner is a semantically reliable and geometry-driven initializer for our semantic-aware parameterization across diverse shapes; and (ii) the visibility-aware loss is required to effectively move seams to less-exposed regions without harming geometric quality. Additional ablation studies are provided in Appendix A.3.

### v3

@@ 3.4 Visibility-Aware UV Parameterization
To address this, we introduce a visibility-aware objective that biases seam placement toward surface areas with low viewer exposure. As a proxy for visibility, we leverage ambient occlusion (AO) (Zhukov et al., 1998), a standard geometric measure of how much of the local hemisphere around a surface point is blocked by nearby geometry (Landis, 2002; Christensen, 2002; Bavoil & Sainz, 2008; McGuire, 2010; Laine & Karras, 2010; Méndez-Feliu & Sbert, 2009). Our training procedure consists of four steps: (i) compute per-vertex AO values on the input mesh, (ii) apply the base UV-parameterization backbone (Sec. 3.2) to obtain candidate UV islands, (iii) extract UV boundary points corresponding to cutting seams, and (iv) compute and minimize the AO-weighted average over the associated 3D vertices. This objective guides the backbone to produce UV islands whose seams coincide with regions of low exposure, thereby relocating cuts to less visible areas of the 3D surface and reducing perceptible seam artifacts after texturing and rendering. We now detail each step in turn.
@@
By this convention, $AO(p)=1$ indicates a fully exposed point (no occlusion) and $AO(p)=0$ indicates a fully occluded point. In practice AO is estimated by Monte Carlo hemisphere sampling (Veach, 1998) or by high-quality offline ray sampling (Pharr et al., 2023) for evaluation. In our implementation, we compute ambient occlusion using the libigl geometry processing library (Jacobson et al., 2018).
Learning Geometry-Preserving Mesh Parameterization. Once per-vertex AO values are computed, we generate the UV parameterization using the base backbone described in Sec. 3.2 and Appendix. A.2 to obtain a potential UV parameterization candidate.
@@
Method Visibility $\downarrow$ Conformality $\uparrow$ Equiareality $\uparrow$ Time (s) $\downarrow$ OptCuts (Li et al., 2018) 0.7855 0.9341 0.8934 240 FlexPara (Single-Chart) (Zhao et al., 2025) 0.8604 0.9097 0.6759 2 Ours - Visibility-Aware Param 0.6065 0.9175 0.6093 2 Ours - Semantic+Visibility-Aware Param 0.6534 0.9153 0.6369 15
Visibility-Aware Seam Loss. After identifying soft seam memberships $s_{i}\in[0,1]$ for each vertex $v_{i}$, we encourage seams to lie in less visible regions of the surface. As a visibility proxy, we use per-vertex ambient occlusion values $AO_{i}\in[0,1]$, computed in the first step, where 0 indicates fully occluded and $1$ indicates fully visible. The loss is defined as the weighted average of occlusion values, using soft seam memberships as weights:
@@
Figure 4: Qualitative results for visibility-aware seam placement and UV parameterization on three representative meshes. For each mesh, the top row shows per-vertex ambient occlusion (yellow = exposed, purple = occluded). Beneath are the visualizations of cutting seams (red) from our method, FlexPara, and OptCuts (top to bottom). Our method places a larger fraction of seam geometry in less-exposed regions, reducing the likelihood of visible seam artifacts under typical viewpoints.
@@ 4 Experiments and Results
Figure 5: Checkerboard texturing using UV maps produced by our visibility-aware method, FlexPara, and OptCuts. Each row shows rendered views of different meshes textured with a checkerboard and a magnified inset of a visually important region near seams (red circles). Because our method steers seams toward occluded regions, the checkerboard pattern appears substantially more continuous from typical camera viewpoints. By contrast, baselines exhibit visible seam artifacts in the zoomed-in insets.
We evaluate our unsupervised semantic- and visibility-aware UV parameterization pipelines on a diverse collection of meshes and compare against several baselines. We provide both qualitative and quantitative comparisons to: FlexPara (Zhao et al., 2025) (recent state-of-the-art global parameterization network), Autodesk Maya (Autodesk, Inc., 2025) (industry-grade commercial software), Blender (Blender Foundation, 2025) (popular open-source modeling tool), and xatlas (jpcy, 2025) (widely-used C++ library for generating unique texture coordinates suitable for lightmap baking and texture painting). Figs. 1,  3,  4, and  5 show representative qualitative results, and Tables. 1 and  3 presents quantitative results of our semantic-aware and visibility-aware pipelines against baselines (additional results are provided in Appendix A.3). In the following, we detail each experiment and its corresponding results.
@@ 4.1 Qualitative Evaluation
Comparative Evaluation of Semantic-Aware UV Parameterization. To evaluate the performance of the proposed semantic-aware UV parameterization pipeline, we compare our method against FlexPara, Autodesk Maya’s Automatic UV generator, Blender’s SmartUV, and xatlas. Fig. 3 shows a representative comparison on a single Bunny mesh; additional results on more diverse and complex meshes are provided in Figs. 11 and 12 in Appendix A.3. In Fig. 3 and Fig. 11, each row corresponds to one method. Semantic parts are color-coded, and columns show the rendered 3D mesh from multiple camera viewpoints. The rightmost column depicts the unwrapped UV coordinates, where UV islands are color-coded consistently with their corresponding semantic 3D parts in the rendered views. As shown, our method successfully identifies semantic parts of the input mesh, parameterizes them, and maps them into UV space with low distortion. This leads to cleaner, more interpretable charts that directly benefit downstream tasks such as texture painting and transfer. By contrast, the baseline methods fail to consistently capture and parameterize semantic parts of the 3D mesh.
Comparative Evaluation of Visibility-Aware UV Parameterization. To evaluate the performance of the proposed visibility-aware UV parameterization pipeline, we compare our method against FlexPara (Zhao et al., 2025) and OptCuts (Li et al., 2018). Figs. 4 and  8 present representative comparisons on different meshes. Additional examples appear in Figs. 5,  9, and 10 in Appendix A.3. In Fig. 4 the top row visualizes per-vertex ambient occlusion (AO): lighter tones (e.g., yellow) indicate high exposure (more visible) while darker tones (e.g., purple) indicate low exposure (more occluded). A desirable parameterization therefore places cutting seams on darker, less-visible regions to reduce perceptible seam artifacts after texturing and rendering. The rows beneath show UV parameterizations produced by our method, FlexPara, and OptCuts (in that order). Qualitatively, our method allocates a larger fraction of seam geometry to low-exposure regions compared to the baselines. Table 1 (and Fig. 7 in Appendix A.3) quantitatively supports this observation: the report the mean AO values of seam vertices for each method, where lower scores indicate seams located in less-visible (more occluded) areas. As shown, our method consistently attains lower mean seam-AO, confirming that the visibility-aware objective successfully biases seam placement toward less-exposed surface regions. Also, Fig. 5 illustrates checkerboard texturing on different mesh objects using parameterizations from our method and from baselines. As shown, our method produces noticeably more seamless texture patterns when rendered from visible viewpoints, whereas FlexPara and OptCuts exhibit prominent cutting seams (highlighted in the zoomed-in insets of Fig. 5).
@@ 4.2 Quantitative Evaluation
User Study. To evaluate semantic- and visibility-awareness of the our method, we conducted a user-preference study with 115 participants. We distributed the questionnaire to two groups: experts and general participants. Of the 115 participants, 70 are general participants (including graduate students with computer science and engineering backgrounds) and 45 are experts: 31 software engineers, 3 project managers, 2 product owners, 5 UV/layout artists, and 4 modeling artists, all working in the Media and Entertainment industry for film and games.
@@
Table 2: To evaluate semantic- and visibility-awareness of the proposed method, we conducted a user‐study with 45 expert participants performing 11 comparisons between textured 3D shapes and UV parameterizations produced by our method and baselines. Our proposed method is strongly preferred by the expert users over the baselines.
@@
Method Expert User Preference Percentage Visibility-Awareness Evaluation OptCuts (Li et al., 2018) 1.48 % FlexPara (single-Chart) (Zhao et al., 2025) 4.81 % Ours - Visibility-Aware Param 93.70 % Semantic-Awareness Evaluation xatlas (jpcy, 2025) 2.66 % Blender (Blender Foundation, 2025) 5.33 % Autodesk Maya (Autodesk, Inc., 2025) 8.44 % FlexPara (multi-Chart) (Zhao et al., 2025) 3.33 % Ours - Semantic-Aware Param 80.22 %
@@
Each participant completed 11 comparisons between textured 3D shapes and UV maps produced by our method, FlexPara, OptCuts, Autodesk Maya, Blender, and xatlas. For each comparison, participants were rated each result according to three visual criteria: (i) texture-pattern smoothness/continuity, (ii) semantic alignment (how well colors correspond to meaningful parts), and (iii) seam visibility (how well seams are placed in occluded/less-obvious regions). The semantic-aware and visibility-aware evaluations comprised 5 and 6 questions, respectively, for a total of 11 comparisons per participant. Table 2 reports the percentage of expert participant preferences for each method. As the table shows, expert users, particularly UV, layout, and modeling artists, strongly prefer our proposed method over the baselines. Examples of questions in our user‐study questionnaire are shown in Fig. 15. Also, Table 10 in Appendix A.3 reports the percentage of general participant preferences for each method. As shown in Table 10, general users also strongly prefer our proposed method over the baselines.
@@
Semantic- and Visibility-Awareness Evaluation. We also quantitatively evaluate both the semantic-aware and visibility-aware components of our pipeline on a collection of representative meshes. For the semantic-aware pipeline we report (i) Hamming distance (lower is better), the fraction of vertices whose labels disagree after optimal label matching to the reference (we use SAMesh (Tang et al., 2024) outputs as a ground-truth); and (ii) Rand Index (RI) (higher is better), a pairwise agreement measure that counts how often vertex pairs are grouped or separated identically between the two labelings. For the visibility-aware pipeline we report Mean seam AO (lower is better in our AO convention, where 0 = occluded and 1 = exposed), the average ambient-occlusion value over seam vertices, which indicates how well seams are placed in occluded regions. Tables 1 and  3 summarize these metrics and compare our method against the baselines described in Sec. 4. Further discussion of the results in these two tables are provided in Appendix A.3.
@@
Geometry-Preservation Evaluation. In addition to task-specific metrics mentioned above, in Tables 1 and  3, we evaluate geometric quality for all methods using conformality (angle-preservation) and equiareality (area-preservation) (higher is better) to ensure the parameterizations remain faithful enough to the original surface. The proposed semantic-aware and visibility-aware UV parameterization frameworks produce an increase in standard distortion metrics including angle-distortion and area-distortion. This is a trade-off which is expected when optimizing multiple objectives during the learning process. Specifically, our optimization introduces explicit semantic-alignment and visibility-awareness terms that encourage semantically consistent charts and reduce conspicuous seams. However, existing geometry-preserving methods (e.g., Autodesk Maya, Blender, xatlas, and FlexPara) focus only on traditional distortion objectives and therefore score better on those numbers, but they do not provide the semantic- and visibility-related guarantees produced by our method.
@@
Method Semantic Awareness Conformality $\uparrow$ Equiareality $\uparrow$ Inference Time (sec) $\downarrow$ Hamming Distance $\downarrow$ Rand Index $\uparrow$ xatlas (jpcy, 2025) 0.8896 0.7023 0.9792 0.9341 12 Blender (Blender Foundation, 2025) 0.8634 0.7173 0.9289 0.9023 << 1 Maya (Autodesk, Inc., 2025) 0.8615 0.7125 0.9272 0.8746 << 1 FlexPara (Multi-Chart) (Zhao et al., 2025) 0.5980 0.6902 0.9592 0.7606 2 Ours - Semantic-Aware Param 0.3188 0.8151 0.9123 0.6707 15 Ours - Semantic+Visibility-Aware Param 0.3212 0.8087 0.9153 0.6369 15
@@
Ablation Studies. To understand which design choices drive our results, we perform a set of controlled ablations and summarize the key findings below (additional qualitative examples are in Fig. 13 and Fig. 14 in Appendix A.3). (i) Partitioning strategy (ShDF vs. SAMesh). We replace our default Shape Diameter Function (ShDF) partitioner (Shapira et al., 2008) with SAMesh (Tang et al., 2024) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively both partitioners produce part decompositions that align with geometric structure across simple and complex models, resulting compact, low-distortion per-part UV charts (see the appendix Fig. 13 and Fig. 14). Quantitatively (Table 6 in Appendix. A.3), using ShDF in our pipeline yields UV parameterizations that are more conformal (better angle preservation) and more equiareal (better area preservation) than those produced when using SAMesh. (ii) Visibility loss (with vs. without). We ablate the proposed visibility-aware objective (Sec. 3.4) by training identical models with and without $\mathcal{L}_{\text{vis}}$. The visibility-aware objective consistently lowers the mean AO of seam vertices (Fig 7 in Appendix. A.3), i.e., it moves seams to regions of lower exposure (lower AO values). Qualitatively, models trained with the AO loss produce seams that are less salient in rendered images from canonical viewpoints (see Fig. 5). Importantly, as presented in Table. 1, adding the visibility term does not noticeably degrade geometric distortion metrics, conformality (angle-preservation) and equiareality area-preservation). The ablations show that: (i) the ShDF partitioner is a semantically reliable and geometry-driven initializer for our semantic-aware parameterization across diverse shapes; and (ii) the visibility-aware loss is required to effectively move seams to less-exposed regions without harming geometric quality. Additional ablation studies are provided in Appendix A.3.